OpenStudy (anonymous):

Is x^2 equivalent to 1/(x^-1)(x^-1)

hartnn (hartnn):

what do u think ?

hartnn (hartnn):

whats \(x^{-1}\times x^{-1} =... \)

OpenStudy (anonymous):

I think it is. The question is to make something equal to x^2. And I made that lol :)

hartnn (hartnn):

usw \(x^m \times x^n=x^{m+n}\)

hartnn (hartnn):

lol, then you are absolutely correct! :) good

OpenStudy (anonymous):

Awesomes! And is \[4\sqrt[4]{x ^{8}} = x ^{2}\]

OpenStudy (anonymous):

Oops, did not mean that coefficient four

hartnn (hartnn):

yes, correct.

hartnn (hartnn):

\(\Large\sqrt[4]{x^8}=x^{8/4}\)

OpenStudy (anonymous):

That was my plan :) Just wanted to make sure I didn't do anything stupid and looked over something

hartnn (hartnn):

you did it correctly :)

OpenStudy (anonymous):

\[x ^{2}= 1/\sqrt[15]{x ^{-30}}\]

OpenStudy (anonymous):

This one I believe is equal as well. Would you say all three I have done so far are different and need at least two steps :)?

hartnn (hartnn):

1st and second combined is 3rd but yes, you can say different

OpenStudy (anonymous):

Awesomes (purposefully misspelled), thank you!

hartnn (hartnn):

lol, welcome ^_^